# Geometric families of multiple elliptic Gamma functions and arithmetic applications, I

### v1

@@ 4.3 Signed fundamental domains
We now prove Theorem 1 using the second approach following [Col88], [DyDF14]. This approach has two benefits: on the one hand the “signed fundamental domain” decomposition described in [DyDF14] is easier to compute than the Shintani decomposition. On the other hand, the cones involved may be interpreted in terms of algebraic cycles (see [CDG15]).
@@
where the $\nu_{\rho}$’s are signs in $\{-1,0,+1\}$ and the $a_{j,\rho}$’s are $\mathbb{Q}$-linear forms on $\mathbb{F}$. To achieve this we will recall the notations from [DyDF14] to describe their signed fundamental domain for the action of $\mathcal{O}^{+,\times}_{\mathfrak{f}}$ on $L=\mathfrak{f}\mathfrak{b}^{-1}$. Fix $\varepsilon_{1},\dots,\varepsilon_{n-1}$ be fundamental units for $\mathcal{O}^{+,\times}_{\mathfrak{f}}$. For any permutation $\rho\in\mathfrak{S}_{n-1}$ and any index $1\leq i\leq n$ define the elements
@@
where the determinants are taken in the canonical basis of $\mathbb{R}^{n}$ and $e_{n}=[0,\dots,0,1]^{t}$ is the last vector of this basis. Then we may define as in [DyDF14] the sets:
@@
Let us then define the cones $C_{\rho}=\sum_{i=1}^{n}\mathbb{R}_{i,\rho}f_{i,\rho}$ for $\rho\in S$ and $C_{\rho}=\sum_{i=1}^{n}\mathbb{R}_{\geq 0}f_{i,\rho}$ if $\rho\not\in S$. We will denote by $c_{\rho}$ the indicator function of the set $C_{\rho}$. It follows from [[DyDF14], Theorem 1] that there are explicit signs $w_{\rho}\in\{-1,0,1\}$ for $\rho\in S$ such that the following equality holds in $\mathcal{K}(V)$:
@@
where $\nu_{\rho}=w_{\rho}(-1)^{\#J_{\rho}}\mathrm{sign}\,\mathrm{det}(a_{1,\rho},\dots,a_{n,\rho})$ if $\rho\in S$ and $\nu_{\rho}=0$ otherwise. Note that the signs $w_{\rho}$ and the cones $C_{\rho}$ appearing in the signed fundamental domain decomposition in [DyDF14] are explicitly computable and so are the linear forms $a_{i,\rho}$ as well as the sets $J_{\rho}$. Using the explicit definition of the $B_{n,a_{1,\rho},\dots,a_{n,\rho}}$ we may rewrite the right-hand side as the trace of an element in $\mathbb{F}$ which implies that $\zeta_{\mathfrak{f}}(\mathfrak{b},0)\in\mathbb{Q}$, as was previously known from the theorem of Klingen and Siegel. $\square$
This second expression of the partial zeta functions using the signed fundamental domain from [DyDF14] is closer to what we had in mind in the discussion carried out in section 3.3 as the cones $C_{\rho}$ admit generators $f_{i,\rho}$ which may be described in terms of an algebraic cycle (see for instance [CDG15]). Yet, we would like the linear forms $a_{i,\rho}$ to be described in terms of similar algebraic cycles. Interesting future work would be to somehow construct a “dual cycle” to the cycle presented in [[Scz93], lemma 5] to be evaluated against the Bernoulli polynomials cocycle. Namely, we wish to construct a linear form $a\in\Lambda$ and define the corresponding cones:
@@
As a last remark on Theorem 1, we stress that it is a reformulation of Shintani’s result, borrowing ideas from [Col88], [DyDF14] and [CDG15], and that there have been many other approaches to the computation of partial zeta values at non-positive integers in totally real fields (see [CGS00] for instance). We argue that the novelty of our work lies in the extraction of arithmetic quantities such as partial zeta values in totally real fields at $s=0$ from the study of the geometric families of $G_{n-2,a_{1},\dots,a_{n-1}}$ functions via the associated geometric families of Bernoulli rational functions $B_{n,a_{1},\dots,a_{n}}$. We will now give two examples of computations of partial zeta values at $s=0$ in real cubic fields following the procedure given in the proof of Theorem 1. All the computations were carried out using the computer software Pari/GP [The23].
@@ 4.4.1 First real cubic example
The corresponding signs $w_{\rho}$ given in [DyDF14] are $w_{\mathrm{Id}}=w_{(12)}=1$. The complete signs $\nu_{\rho}$ are $\nu_{\mathrm{Id}}=-1$ and $\nu_{(12)}=1$. Following the proof of Theorem 1 we set $a_{i,\mathrm{Id}}=7\mu_{i,\mathrm{Id}}b_{i,\mathrm{Id}}$ and $a_{i,(12)}=97\mu_{i,(12)}b_{i,(12)}$ for $1\leq i\leq 3$. We may then compute using formula (14):

### v2

@@ 4.3 Signed fundamental domains
We now prove Theorem 1 using the second approach following [COL88], [DF14]. This approach has two benefits: on the one hand the “signed fundamental domain” decomposition described in [DF14] is easier to compute than the Shintani decomposition. On the other hand, the cones involved may be interpreted in terms of algebraic cycles (see [CDG15]).
@@
where the $\nu_{\rho}$’s are signs in $\{-1,0,+1\}$ and the $a_{j,\rho}$’s are $\mathbb{Q}$-linear forms on $\mathbb{F}$. To achieve this we will recall the notations from [DF14] to describe their signed fundamental domain for the action of $\mathcal{O}^{+,\times}_{\mathfrak{f}}$ on $L=\mathfrak{f}\mathfrak{b}^{-1}$. Fix $\varepsilon_{1},\dots,\varepsilon_{n-1}$ be fundamental units for $\mathcal{O}^{+,\times}_{\mathfrak{f}}$. For any permutation $\rho\in\mathfrak{S}_{n-1}$ and any index $1\leq i\leq n$ define the elements
@@
where the determinants are taken in the canonical basis of $\mathbb{R}^{n}$ and $e_{n}=[0,\dots,0,1]^{t}$ is the last vector of this basis. Then we may define as in [DF14] the sets:
@@
Let us then define the cones $C_{\rho}=\sum_{i=1}^{n}\mathbb{R}_{i,\rho}f_{i,\rho}$ for $\rho\in S$ and $C_{\rho}=\sum_{i=1}^{n}\mathbb{R}_{\geq 0}f_{i,\rho}$ if $\rho\not\in S$. We will denote by $c_{\rho}$ the indicator function of the set $C_{\rho}$. It follows from [[DF14], Theorem 1] that there are explicit signs $w_{\rho}\in\{-1,0,1\}$ for $\rho\in S$ such that the following equality holds in $\mathcal{K}(V)$:
@@
where $\nu_{\rho}=w_{\rho}(-1)^{\#J_{\rho}}\mathrm{sign}\,\mathrm{det}(a_{1,\rho},\dots,a_{n,\rho})$ if $\rho\in S$ and $\nu_{\rho}=0$ otherwise. Note that the signs $w_{\rho}$ and the cones $C_{\rho}$ appearing in the signed fundamental domain decomposition in [DF14] are explicitly computable and so are the linear forms $a_{i,\rho}$ as well as the sets $J_{\rho}$. Using the explicit definition of the $B_{n,a_{1,\rho},\dots,a_{n,\rho}}$ we may rewrite the right-hand side as the trace of an element in $\mathbb{F}$ which implies that $\zeta_{\mathfrak{f}}(\mathfrak{b},0)\in\mathbb{Q}$, as was previously known from the theorem of Klingen and Siegel. $\square$
This second expression of the partial zeta functions using the signed fundamental domain from [DF14] is closer to what we had in mind in the discussion carried out in section 3.3 as the cones $C_{\rho}$ admit generators $f_{i,\rho}$ which may be described in terms of an algebraic cycle (see for instance [CDG15]). Yet, we would like the linear forms $a_{i,\rho}$ to be described in terms of similar algebraic cycles. Interesting future work would be to somehow construct a “dual cycle” to the cycle presented in [[SCZ93], lemma 5] to be evaluated against the Bernoulli polynomials cocycle. Namely, we wish to construct a linear form $a\in\Lambda$ and define the corresponding cones:
@@
As a last remark on Theorem 1, we stress that it is a reformulation of Shintani’s result, borrowing ideas from [COL88], [DF14] and [CDG15], and that there have been many other approaches to the computation of partial zeta values at non-positive integers in totally real fields (see [BEK24] or [CGS00] for instance). We argue that the novelty of our work lies in the extraction of arithmetic quantities such as partial zeta values in totally real fields at $s=0$ from the study of the geometric families of $G_{n-2,a_{1},\dots,a_{n-1}}$ functions via the associated geometric families of Bernoulli rational functions $B_{n,a_{1},\dots,a_{n}}$. We will now give two examples of computations of partial zeta values at $s=0$ in real cubic fields following the procedure given in the proof of Theorem 1. All the computations were carried out using the computer software Pari/GP [18].
@@ 4.4.1 First real cubic example
The corresponding signs $w_{\rho}$ given in [DF14] are $w_{\mathrm{Id}}=w_{(12)}=1$. The complete signs $\nu_{\rho}$ are $\nu_{\mathrm{Id}}=-1$ and $\nu_{(12)}=1$. Following the proof of Theorem 1 we set $a_{i,\mathrm{Id}}=7\mu_{i,\mathrm{Id}}b_{i,\mathrm{Id}}$ and $a_{i,(12)}=97\mu_{i,(12)}b_{i,(12)}$ for $1\leq i\leq 3$. We may then compute using formula (14):